OpenStudy (anonymous):

Find DxY using the rules for finding derivatives: y=(3x^2-5x+7)/(5x+9)

OpenStudy (shamil98):

Use the quotient rule.

OpenStudy (shamil98):

f(x) = 3x^2-5x+7 f'(x) = ? g(x) = 5x + 9 g'(x) =? From there just input those values into the quotient rule: \[y' = \frac{ f'(x)g(x) - f(x)g'(x) }{(g(x))^2}\]